Suppose the national reading level for high school students is 140 words per minute with a standard deviation of 12. A local high school wants to know if their students read at a level different from the national average. The level of the test is set at .20. A random sample size of 125 students has been drawn and the resulting average is 132 words per minute. Provide the critical value, the test statistic and the number of tails this hypothesis test has.
Answer #1

Solution :

Given that ,

\mu = 140

T = 132

\sigma = 12

n = 125

The null and alternative hypothesis is ,

H0 :  \mu = 140

Ha : \mu   \neq 140

This is the two tailed test .

\alpha = 0.20

Z\alpha = Z0.20 = +/-1.28

The critical value = +/-1.28

Test statistic = z

= (T - \mu ) / \sigma / \sqrt n

= ( 132 -140) / 12 / \sqrt 125

= -7.45

The test statistic = -7.45